Find the next three terms 0, 3, 8, 15, 24
Answer by Edwin McCravy(20056)

 0 +  3 =  3
 3 +  5 =  8
 8 +  7 = 15
15 +  9 = 24
24 + 11 = 35
35 + 13 = 48
48 + 15 = 63

Edwin